5. Determine the length of XY to the nearest tenth of a centimetre.


61°

17.4 cm

Y X


Z

a. 8.4 cm b. 15.2 cm c. 31.4 cm d. 19.9 cm


1
Expert's answer
2021-05-30T23:16:17-0400

To get angle XZY="90-61=29\u00b0"


Using the sine rule ;

"\frac {17.4} {Sin 90}=\frac {x} {Sin 29}"


"\frac {17.4} {1}=\frac {x} {0.4848}"


"x=(17.4\u00d70.4848)=8.4cm"


XY=8.4cm


a) 8.4 cm Correct